### v4

@@ 2 Background
[…]
@@
[…]
@@ 3.1 Problem Definition
[…]
@@ 3.3 Theory
where $Y_{n,k}$ is the empirical distribution of a replacement sample with at least $n-k$ points from the original sample $X$. That is, $Y_{n,k}$ contains at least $n-k$ observations drawn without replacement from $X$ and the rest consists of arbitrary points from $M$. The finite-sample breakdown point for an unbounded Riemannian manifold was examined in Fletcher et al. (2009). This result allows direct extension to the 2-Wasserstein space, which we introduce for completeness.
@@ 4.2 Gaussian Distributions
0:  $\Sigma^{*}=\operatorname*{argmin}_{\Sigma\in\mathcal{N}_{0}^{d}}\sum_{n=1}^{N}\pi_{n}W_{2}(\Sigma,\Sigma_{n})$.
@@ 5.1 Simulated Example
Our experiment procedure is as follows. It starts by generating perturbed versions of a distribution by drawing a random sample of varying size $n$ from each distribution, from which an empirical cumulative distribution function (CDF) is computed. This procedure is repeated $(100-k)$ times for Gamma(1,1) and $k$ times for Gamma(7.5, 0.75), leading to a collection of 100 empirical distributions. We set the number of sampling from type 2 measure as $k\in[1,25]$. Since we consider $k$ empirical distributions as contamination out of a total of 100 distributions, we will denote the degree of contamination as $k\%$. When estimates of the Wasserstein barycenter and median are computed from the collection of 100 empirical CDFs are obtained, the discrepancy between an estimate and the signal - Gamma(1,1) distribution - is measured by the 2-Wasserstein distance. This procedure elaborates how much of an error is induced for a centroid estimator in the presence of outliers.
@@ 5.2 Real Example 1 : Newcomb’s Speed of Light
The first real data example involves Newcomb’s measurements of the speed of light (Newcomb, 1891). The data are composed of 66 scalar-valued measurements collected on different days recorded in millionths of a second. In a version of the dataset where the true speed of light is scaled to $33$, out of 66 observations, two measurements are negative: $-2$ and $-44$. Putting the possibility of time travel aside, it is reasonable to attribute this phenomenon to measurement errors and view these observations as outliers.
@@ 5.3 Real Example 2 : MNIST Digits
Figure 7: Visualization of various centroids from the MNIST example. Each row represents the Euclidean mean (row 1), the Euclidean median (row 2), the Wasserstein barycenter (3rd row), and the Wasserstein median (4th row) of 50 images per digit.
@@ A. Proof of Theoretical Results
[…]
@@
[…]
@@
[…]
@@ B.1. Centered Gaussian Measures
As shown above, an identity matrix ${\bf\Sigma}_{1}$ is drawn as a circle of radius 1 and ${\bf\Sigma}_{2}$ is a rotated ellipse. Similar to the univariate distribution example from the manuscript, we generate perturbed variants of each distribution by computing a sample covariance for a randomly generated sample from the distribution. We repeat this process $(100-k)$ times for $N({\bf\Sigma}_{1})$ and $k$ times for $N({\bf\Sigma}_{2})$ for $k\in[1,25]$. This simulates a scenario where a majority of centered Gaussian measures resembles the signal measure $N({\bf\Sigma}_{1})$ and a small portion of perturbation comes from the contamination $N({\bf\Sigma}_{2})$. When barycenter and median estimates are obtained, we report the discrepancy between the estimates and the signal measure $N({\bf\Sigma}_{1})$ by the 2-Wasserstein distance.
@@ B.2. High-dimensional Gaussian Measures
As before, we generated a random sample of size $2p$ from $N({\bf\Sigma_{1}})$ and estimated its empirical covariance, resulting in a perturbed realization of ${\bf\Sigma_{1}}$. This sampling and estimation process was repeated $100-k$ times. Similarly, we repeated this process $k$ times for ${\bf\Sigma_{2}}$. Aggregating these two sets of covariances yielded a collection of 100 covariance matrices, with $k$ representing contaminations. Then, the Wasserstein barycenter and median were computed according to the Bures-Wasserstein geometry of centered Gaussian distributions, which we denote as $\hat{\Sigma}_{\text{mean}}$ and $\hat{\Sigma}_{\text{med}}$, respectively. Our experiment considers different degrees of contamination for $k=1,5,10,25$ across varying dimensions $p=16,32,\ldots,1024$. Also, we added another dimension of comparison for $\rho=0.1,0.5,0.9$. For each setting, we quantified the comparative performance using the following metric
@@
Noise Level Dimensionality
@@
Noise Level Dimensionality
@@
Noise Level Dimensionality
@@ B.3. Histograms
Histogram is another popular representation of the distribution, which can be considered as a discrete measure supported on one-dimensional lattice by considering the midpoints of each bin as points of the lattice. Similar to before, we consider a signal-noise setting to see applicability of the proposed framework on a collection of histogram objects. In this example, we opt for the beta distribution denoted as Beta($\alpha,\beta$) for two shape parameters $\alpha>0$ and $\beta>0$ as a model distribution. We consider Beta(2,5) as the signal measure whose density function has a mound shape and positive skewness. For the contamination measure, we use Beta(5,1) whose density function is monotonically increasing in the bounded support. We use an equidistant partition $10=t_{0}<\cdots<t_{20}=$ 1 for binning a randomly generated sample and normalize bin counts as relative frequency so that the binned vector sums to 1.

### v5

@@ 2 Background
$\mathcal{P}_{p}(\mathbb{R}^{d})=\left\{\mu\in\mathcal{P}(\mathbb{R}^{d}):\int_{\mathbb{R}^{d}}\|\mathbf{x}\|^{p}\,d\mu(\mathbf{x})<\infty\right\},\quad p\geq 1,$
where $\|\cdot\|$ is the standard norm in the Euclidean space. The distance for any $\mu,\nu\in\mathcal{P}_{p}(\mathbb{R}^{d})$ is defined as the minimum of total transportation cost
$\mathcal{W}_{p}(\mu,\nu)=\left(\underset{T:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}}{\inf}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\|{\bf x}-T({\bf x})\|^{p}\,d\mu({\bf x})\right)^{1/p},$ (1)
@@
$\mathcal{W}_{p}(\mu,\nu)=\left(\underset{\pi\in\Pi(\mu,\nu)}{\inf}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\|{\bf x}-{\bf y}\|^{p}\,d\pi({\bf x},{\bf y})\right)^{1/p},$ (2)
@@ 3.1 Problem Definition
$F(\nu)=\mathbb{E}\mathcal{W}_{2}(\nu,\Lambda)=\int\mathcal{W}_{2}(\nu,\lambda)d\mathbb{P}(\lambda),$ (4)
@@ 3.3 Theory
where $Y_{n,k}$ is the empirical distribution of a replacement sample with at least nkn-k points from the original sample $X$. That is, $Y_{n,k}$ contains at least nkn-k observations drawn without replacement from $X$ and the rest consists of arbitrary points from $M$. The finite-sample breakdown point for an unbounded Riemannian manifold was examined in Fletcher et al. (2009). This result allows direct extension to the 2-Wasserstein space, which we introduce for completeness.
@@ 4.2 Gaussian Distributions
[…]
@@ 5.1 Simulated Example
Our experiment procedure is as follows. It starts by generating perturbed versions of a distribution by drawing a random sample of varying size $n$ from each distribution, from which an empirical cumulative distribution function (CDF) is computed. This procedure is repeated (100k)(100-k) times for Gamma(1,1) and $k$ times for Gamma(7.5, 0.75), leading to a collection of 100 empirical distributions. We set the number of sampling from type 2 measure as $k\in[1,25]$. Since we consider $k$ empirical distributions as contamination out of a total of 100 distributions, we will denote the degree of contamination as $k\%$. When estimates of the Wasserstein barycenter and median are computed from the collection of 100 empirical CDFs are obtained, the discrepancy between an estimate and the signal - Gamma(1,1) distribution - is measured by the 2-Wasserstein distance. This procedure elaborates how much of an error is induced for a centroid estimator in the presence of outliers.
@@ 5.2 Real Example 1 : Newcomb’s Speed of Light
The first real data example involves Newcomb’s measurements of the speed of light (Newcomb, 1891). The data are composed of 66 scalar-valued measurements collected on different days recorded in millionths of a second. In a version of the dataset where the true speed of light is scaled to $33$, out of 66 observations, two measurements are negative: 2-2 and 44-44. Putting the possibility of time travel aside, it is reasonable to attribute this phenomenon to measurement errors and view these observations as outliers.
@@ 5.3 Real Example 2 : MNIST Digits
Figure 7: Visualization of various centroids from the MNIST example. Each row represents the Euclidean mean (1st row), the Euclidean median (2nd row), the Wasserstein barycenter (3rd row), and the Wasserstein median (4th row) of 50 images per digit.
@@ A. Proof of Theoretical Results
If the functional $F(\nu)=\mathbb{E}\mathcal{W}_{2}(\nu,\Lambda)$ is identically infinite, the statement is valid for any choice of measure in $\mathcal{P}_{2}(\mathbb{R}^{d})$. If not, take a minimizing sequence $\{\nu_{j}\}\subset\mathcal{P}_{2}(\mathbb{R}^{d})$ and denote an upper bound of the functional with respect to the sequence as $C=\underset{j}{\sup}~F(\nu_{j})<\infty$. Assume $\nu_{0}$ be a fixed reference measure. By the triangle inequality, for a Dirac measure $\delta_{0}$ in $\mathcal{P}_{2}(\mathbb{R}^{d})$, we have
@@
$\nu_{j}\left(\left\{{\bf x}\in\mathbb{R}^{d}~|~\|{\bf x}\|\geq R\right\}\right)\leq\frac{1}{R^{2}}\int_{\{{\bf x}:\|{\bf x}\|\geq R\}}\|{\bf x}\|^{2}d\nu_{n}({\bf x})\leq\left(\frac{M}{R}\right)^{2}.$
@@
$\geq\mathbb{E}\liminf_{j_{k}\rightarrow\infty}\mathcal{W}_{2}(\nu_{j_{k}},\mu_{j_{k}})\text{ by Fatou's lemma}$
@@ B.1. Centered Gaussian Measures
As shown above, an identity matrix ${\bf\Sigma}_{1}$ is drawn as a circle of radius 1 and ${\bf\Sigma}_{2}$ is a rotated ellipse. Similar to the univariate distribution example from the manuscript, we generate perturbed variants of each distribution by computing a sample covariance for a randomly generated sample from the distribution. We repeat this process (100k)(100-k) times for $N({\bf\Sigma}_{1})$ and $k$ times for $N({\bf\Sigma}_{2})$ for $k\in[1,25]$. This simulates a scenario where a majority of centered Gaussian measures resembles the signal measure $N({\bf\Sigma}_{1})$ and a small portion of perturbation comes from the contamination $N({\bf\Sigma}_{2})$. When barycenter and median estimates are obtained, we report the discrepancy between the estimates and the signal measure $N({\bf\Sigma}_{1})$ by the 2-Wasserstein distance.
@@ B.2. High-dimensional Gaussian Measures
As before, we generated a random sample of size $2p$ from $N({\bf\Sigma_{1}})$ and estimated its empirical covariance, resulting in a perturbed realization of $\bf\Sigma_{1}$}. This sampling and estimation process was repeated 100k100-k times. Similarly, we repeated this process $k$ times for $\bf\Sigma_{2}$}. Aggregating these two sets of covariances yielded a collection of 100 covariance matrices, with $k$ representing contaminations. Then, the Wasserstein barycenter and median were computed according to the Bures-Wasserstein geometry of centered Gaussian distributions, which we denote as $\hat{\Sigma}_{\text{mean}}$ and $\hat{\Sigma}_{\text{med}}$, respectively. Our experiment considers different degrees of contamination for $k=1,5,10,25$ across varying dimensions $p=16,32,\ldots,1024$. Also, we added another dimension of comparison for $\rho=0.1,0.5,0.9$. For each setting, we quantified the comparative performance using the following metric
@@
Noise
Level
Dimensionality
@@
Noise
Level
Dimensionality
@@
Noise
Level
Dimensionality
@@ B.3. Histograms
Histogram is another popular representation of the distribution, which can be considered as a discrete measure supported on one-dimensional lattice by considering the midpoints of each bin as points of the lattice. Similar to before, we consider a signal-noise setting to see applicability of the proposed framework on a collection of histogram objects. In this example, we opt for the beta distribution denoted as Beta($\alpha,\beta$) for two shape parameters $\alpha>0$ and $\beta>0$ as a model distribution. We consider Beta(2,5) as the signal measure whose density function has a mound shape and positive skewness. For the contamination measure, we use Beta(5,1) whose density function is monotonically increasing in the bounded support. We use an equidistant partition $0=t_{0}<\cdots<t_{20}=1$ for binning a randomly generated sample and normalize bin counts as relative frequency so that the binned vector sums to 1.